4. A box contains 500 envelopes, of which 75 contains P5,000 in cash, 150 contains P1,250, and 275 contain P500. An envelope may be purchased for P1,250. What is the sample space for the different amounts of money? Assign probabilities to the sample points and then find the probability that the first envelope purchased contains less than P5,000.
a. P(5000)
b. P(1250)
c. P(500)
d. P(x<5000)
1
Expert's answer
2020-10-20T18:49:20-0400

S={500,1250,5000}

a)P(x=5000)=75/500=0.15

b)P(x= 1250)=150/500=0.3

c)P(x=500)=275/500=0.55

d)P(x<5000)=P(x=500)+P(x=1250)=0.55+0.3=0.85